OpenStudy (anonymous):

Which of the following is the correct factored form of the given equation? x2 + 3x - 4 = 0 (x - 1)(x + 4) = 0 (x - 1)(x - 4) = 0 (x + 1)(x - 4) = 0

OpenStudy (anonymous):

\[ x^2 + 3x - 4 = x^2 + 4x - x - 4 = (x-1)(x-4) =0 \]

OpenStudy (kirbykirby):

No it's (x-1)(x+4)

OpenStudy (kirbykirby):

Which two numbers allow you to get 4 when multiplied, but 3 when added together? -1 and 4, hence (x - 1)(x + 4) If you expand it you get x^2 + 4x - x - 4 = x^2 +3x -4

OpenStudy (anonymous):

think about which factors of 4 (1,2,and,4) will add or subtract to equal 3. it must by 4 minus 1 = 3

OpenStudy (anonymous):

which is it ?

OpenStudy (kirbykirby):

If you need more proof, here's a link http://www.wolframalpha.com/input/?i=x^2%2B3x-4%3D0+factor